# Invariant measures for some dissipative systems from the Jacobi last multiplier

### v1

@@ 3.2 Contact Hamiltonian systems
In this section, let us discuss the role that the Jacobi last multiplier plays in the context of contact Hamiltonian systems in describing invariant phase-space measures. For a generic contact Hamiltonian system, i.e., the triple $(\mathcal{M}_{c},\eta,h)$, equations (2.2) and (2.25) imply that
$\frac{d}{dt}\ln M=(n+1)\xi(h).$ (3.8)
Let us consider two distinct cases below in which we will re-derive the results of [11] as suited for contact Hamiltonian systems but using the framework of the last multiplier.
@@ 3.2.1 Level set $\{h^{-1}(0)\}$
In the region of the (contact) phase space where $h=0$, (2.23) implies that the flow of the contact vector field $X_{h}$ should be confined to within that region as $h$ is conserved when $h=0$. In other words, the level set $\{h^{-1}(0)\}$ is invariant under the contact Hamiltonian dynamics. This implies the following result:
Theorem 3.2
On the level set $\{h^{-1}(0)\}$, the Jacobi last multiplier for a contact Hamiltonian system $(\mathcal{M}_{c},\eta,h)$ is a differentiable function of $h$, i.e.,
$M=M(h).$ (3.9)
Corollary 3.2
On the level set $\{h^{-1}(0)\}$, the following is an invariant measure:
$\Omega\big{|}_{h=0}=\big{[}M(h)(\eta\wedge(d\eta)^{n})\big{]}_{h=0},$ (3.10)
Remark: Notice that the last multiplier is not unique but assumes a constant value on the level set $\{h^{-1}(0)\}$.
@@ 3.2.2 Region $\mathcal{M}_{c}\setminus\{h^{-1}(0)\}$
Consider a contact Hamiltonian system $(\mathcal{M}_{c},\eta,h)$ where ${\rm dim~{}}\mathcal{M}_{c}=2n+1$. In the region $\mathcal{M}_{c}\setminus\{h^{-1}(0)\}$, the Jacobi last multiplier is given by
@@ 4.1 Generalized conformal vector fields
Consider a cotangent bundle $\mathcal{M}_{s}=T^{*}Q$ which is naturally equipped with the tautological one-form $\theta$ leading to the symplectic two-form $\omega=d\theta$. Given a function $K\in C^{\infty}(Q,\mathbb{R})$, we can define a generalized conformal vector field $X^{K}_{H}$ as
£$\pounds_{X^{K}_{H}}\omega=-K\omega,$ (4.1)
and which is compatible with the condition $\iota_{X^{K}_{H}}\omega=dH-K\theta$ because $dK\wedge\theta=0$.
In Darboux coordinates $(q^{i},p_{i})$, we have $K=K(q^{i})$ and the vector field $X^{K}_{H}$ reads as
$X^{K}_{H}=\frac{\partial H}{\partial p_{i}}\frac{\partial}{\partial q^{i}}-\bigg{(}\frac{\partial H}{\partial q^{i}}+K(q^{i})p_{i}\bigg{)}\frac{\partial}{\partial p_{i}}.$ (4.2)
For simplicity, consider a two-dimensional phase space with a symplectic form $\omega=dq\wedge dp$. For some $H=H(q,p)$ and $K=K(q)$, the equations of motion turn out to be
@@
Interpreting $q$ as a mechanical coordinate with $p$ being the corresponding momentum, one finds that the second equation above describes dynamics with dissipation linear in the momentum but with a position-dependent damping strength. When $K$ is a constant, the dynamics described above reduces to that dictated by a conformal vector field (2.13) discussed earlier.
Definition 4.2
Consider an exact symplectic manifold $(\mathcal{M}_{s},\omega)$ which is a cotangent bundle $T^{*}Q$ with the tautological one-form $\theta$, i.e., $\omega=d\theta$. Then, a generalized conformal Hamiltonian system is the quadruple $(\mathcal{M}_{s},\omega,H,K)$, where $H\in C^{\infty}(\mathcal{M}_{s},\mathbb{R})$ and $K\in C^{\infty}(Q,\mathbb{R})$. The corresponding dynamics is described by a generalized conformal vector field which is determined through the condition $\iota_{{X}^{K}_{H}}\omega=dH-K\theta$.
Using the Liouville vector field $\Delta$ on a cotangent bundle as defined in (2.16), one can express a generalized conformal vector field as
@@
where $X_{H}$ is the (conservative) Hamiltonian vector field corresponding to the function $H\in C^{\infty}(\mathcal{M}_{s},\mathbb{R})$. Thus, it turns out that the Hamiltonian is not conserved, i.e.,
@@
and from (4.1), the divergence of $X^{K}_{H}$ turns out to be ${\rm div}\cdot X^{K}_{H}=-nK$, where $\omega^{n}$ is the chosen volume-form on $(\mathcal{M}_{s},\omega)$ which is not invariant under the flow of $X^{K}_{H}$.
@@ 4.3 Cheillini integrability condition and invariant measures
Now that we have established the correspondence between generalized conformal Hamiltonian dynamics and Liénard-type systems, let us proceed towards deriving an expression for the Jacobi last multiplier which will allow us to describe invariant phase-space measures. We will prove the following result:
@@
where $l\neq-1$ and is determined from the condition (4.9). If $K(q)$ and $V(q)$ are such that (4.9) is satisfied, then we get $\dot{u}=luK(q)$ along with $p=mu+G(q)$ which is the same as the Liénard system. However, now that one has $\dot{u}=luK(q)$, we can write $K(q)dt=(lu)^{-1}du$ and substituting this into (4.13) gives us $M=u^{1/l}$ or equivalently,
@@
Remarks: The expressions (4.10) and (4.15) are true only when the condition (4.9) is true; (4.9) is termed as the Cheillini integrability condition (see for example, [8, 10, 21, 22]). Notice that generally two values of $l$ are obtained from (4.9) and therefore one obtains two invariant measures from (4.10). Thus, the invariant measures are not unique.
5 Closing remarks
In this paper, we have presented invariant phase-space measures of certain dissipative systems using the framework of the Jacobi last multiplier. In particular, we analyzed conformal Hamiltonian dynamics as well as contact Hamiltonian dynamics; for the latter, the results of [11] were obtained using the framework of the last multiplier. Another class of systems that we analyzed were systems of the Liénard class which may possess position-dependent damping. For such systems, we discussed a geometric description of the dynamics on cotangent bundles by presenting the notion of a generalized conformal vector field. This also allowed us to obtain analytical expressions for invariant phase-space measures associated with such systems when the Cheillini integrability condition is satisfied. A closely-related analysis may be carried out for a class of Levinson-Smith equations discussed recently in [22].

### v2

@@ 3.2 Contact Hamiltonian systems
In this section, let us discuss the role that Jacobi last multipliers play in the context of contact Hamiltonian systems in describing invariant phase-space measures. For a generic contact Hamiltonian system, i.e., the triple $(\mathcal{M}_{c},\eta,h)$, equations (2.2) and (2.25) imply that
$X_{h}(\ln M)=(n+1)\xi(h).$ (3.8)
Let us consider two distinct cases below. In particular, in Sec. (3.2.2), we will re-derive the ‘canonical invariant measure’ as suited for contact Hamiltonian systems (presented originally in [11]) using the framework of last multipliers.
@@ 3.2.1 Level set $\{h^{-1}(0)\}$
[…]
@@ 3.2.2 Region $\mathcal{M}_{c}\setminus\{h^{-1}(0)\}$
Consider a contact Hamiltonian system $(\mathcal{M}_{c},\eta,h)$ where ${\rm dim~{}}\mathcal{M}_{c}=2n+1$. In the region $\mathcal{M}_{c}\setminus\{h^{-1}(0)\}$, there is a Jacobi last multiplier given by
@@
3.2.3 Constant-damping systems
For a linearly-damped system with constant damping as treated in Sec. (2.4.2), it is easy to apply the result (3.12). On the three-dimensional phase space with $\eta\wedge d\eta=ds\wedge dq\wedge dp$, the invariant phase-space measure reads
$\Omega|_{h\neq 0}=\bigg{[}\frac{p^{2}}{2m}+V(q)+\gamma s\bigg{]}^{-2}(ds\wedge dq\wedge dp),\quad\quad h=\frac{p^{2}}{2m}+V(q)+\gamma s\neq 0.$ (3.13)
@@ 4.1 Generalized conformal vector fields
Consider a configuration space $Q$ and a corresponding phase space $T^{*}Q$ endowed with a symplectic two-form $\omega$. A generalized conformal vector field $Y$ is a vector field such that there exists a function $K$ on $Q$ such that
£$\pounds_{Y}\omega=-K\omega,$ (4.1)
and then there exists a (locally defined) function $H$ in $T^{*}Q$ such that $\iota_{Y}\omega=dH-K\theta$. Such a vector field $Y$ will be denoted by $X^{K}_{H}$.
Corollary 4.1
Consider the cotangent bundle $T^{*}\mathbb{R}$ which is naturally equipped with the tautological one-form $\theta$ leading to the symplectic two-form $\omega=d\theta$. If $X^{K}_{H}$ be a generalized conformal vector field and if the function $K$ defined on $\mathbb{R}$ is differentiable, then the condition £$\pounds_{X^{K}_{H}}\omega=-K\omega$ is equivalent to $\iota_{X^{K}_{H}}\omega=dH-K\theta$ because $dK\wedge\theta=0$.
Proof – Consider the two-dimensional phase space $\pi:T^{*}\mathbb{R}\rightarrow\mathbb{R}$. If $q$ be a coordinate on $\mathbb{R}$ and $\pi:(q,p)\rightarrow q$ be the corresponding ‘induced’ fiber coordinates on the cotangent bundle, then naturally, $\theta=-pdq$ (implying $\omega=d\theta=dq\wedge dp$) from the coordinate definition of the tautological one-form on $T^{*}\mathbb{R}$. Thus, given a differentiable function $K(q)$ on $\mathbb{R}$, one has $dK\wedge\theta=0$. Consequently, from Cartan’s formula (2.4) for the Lie derivative, one can write £$\pounds_{X^{K}_{H}}\omega=d(\iota_{X^{K}_{H}}\omega)+\iota_{X^{K}_{H}}(d\omega)$, implying that $-K\omega=d(\iota_{X^{K}_{H}}\omega)$. The two sides of this equation agree if we put $\iota_{X^{K}_{H}}\omega=dH-K\theta$ because $dK\wedge\theta=0$.
Since our interest is in Liénard-type systems which admit one-dimensional configuration spaces, we will focus hereafter on phase spaces that are $T^{*}\mathbb{R}$ (or a subspace of it) for which the above-mentioned corollary holds. For some $H=H(q,p)$ and $K=K(q)$, the equations of motion turn out to be
@@
Interpreting $q$ as a mechanical coordinate with $p$ being the corresponding momentum, one finds that the second equation above describes dynamics with damping linear in the momentum but with a position-dependent damping strength. When $K$ is a constant, the dynamics described above reduces to that dictated by a conformal vector field (2.13) discussed earlier.
Using the Liouville vector field $\Delta$ on $T^{*}\mathbb{R}$ as defined in (2.16), one can express a generalized conformal vector field as
@@
where $X_{H}$ is the (conservative) Hamiltonian vector field corresponding to the function $H\in C^{\infty}(T^{*}\mathbb{R},\mathbb{R})$. Thus, it turns out that the Hamiltonian is not conserved, i.e.,
@@
and from (4.1), the divergence of $X^{K}_{H}$ turns out to be ${\rm div}X^{K}_{H}=-K$, where $\omega$ is the chosen volume-form on $(\mathcal{M}_{s}=T^{*}\mathbb{R},\omega)$ which is not invariant under the flow of $X^{K}_{H}$.
@@ 4.3 Cheillini integrability condition and invariant measures
We will prove the following result:
@@
where $l$ is determined from the condition (4.8). If $K(q)$ and $V(q)$ are such that (4.8) is satisfied, we then get $\dot{u}=luK(q)$ along with $p=mu+G(q)$ which is equivalent to the Liénard system. However, now that one has $\dot{u}=luK(q)$, we can write $K(q)dt=(lu)^{-1}du$ and substituting this into (4.12) gives us $M=u^{1/l}$ or equivalently,
@@
Remarks: The expressions (4.9) and (4.14) are true only when the condition (4.8) is true; (4.8) is equivalent to (1.2) which is termed the Cheillini integrability condition (see for example, [8, 10, 22, 23]). Notice that generally two values of $l$ are obtained from (4.8) and therefore the invariant measures are not unique.
4.3.1 Constant-damping systems in quadratic and constant potentials
If one considers the case of constant damping as a special case of the Liénard system, then putting $f(q)=K(q)=\gamma>0$ implies that the Cheillini integrability condition (4.8) demands
$V^{\prime\prime}(q)+m\gamma^{2}l(l+1)=0.$ (4.15)
Let us take two cases one after the other.
Quadratic potential: Choosing $V(q)=\alpha_{0}q^{2}$ where $\alpha_{0}$ is a constant, one gets the quadratic equation
$l^{2}+l+\frac{2\alpha_{0}}{m\gamma^{2}}=0.$ (4.16)
This generally gives two values for $l$, whose reality imposes the condition $m\gamma^{2}\geq 8\alpha_{0}$. For simplicity, choosing $\alpha_{0}=m\gamma^{2}/8$ so that there is only one value of $l$, namely, $l=-1/2$, the expression (4.14) gives
$M(q,p)=\bigg{[}p+\frac{m\gamma q}{2}\bigg{]}^{-2},$ (4.17)
up to a constant factor. Using this, it is easy to check that $X^{K=\gamma}_{H}(\ln M)=\gamma$, which satisfies the equation (2.2). Thus, the invariant measure (4.9) is
$\Omega=\bigg{[}p+\frac{m\gamma q}{2}\bigg{]}^{-2}(dq\wedge dp),$ (4.18)
subject to $p\neq\frac{m\gamma q}{2}$. Notice that we cannot treat general potentials here as they will not satisfy the Cheillini integrability condition (4.8).
Constant potential: If $V^{\prime}(q)=0$, the Cheillini integrability condition (4.8) implies $l=0,-1$. Referring to equation (4.14), one must discard $l=0$ and take $l=-1$, giving $M\sim 1/p$. Thus, the result (4.9) now corresponds to the result (3.7) obtained earlier for the free particle with constant damping.
5 Discussion
In this paper, we have presented invariant phase-space measures for certain mechanical systems using the framework of Jacobi last multipliers. In particular, we analyzed conformal Hamiltonian dynamics as well as contact Hamiltonian dynamics. Another class of systems that we analyzed were systems of the Liénard class which may possess position-dependent damping. For such systems, we discussed a geometric description of the dynamics on cotangent bundles by presenting the notion of a generalized conformal vector field. This also allowed us to obtain analytical expressions for invariant phase-space measures associated with such systems when the Cheillini integrability condition is satisfied. A closely-related analysis may be carried out for a class of Levinson-Smith equations discussed recently in [23].
Let us end by interpreting the notion of last multipliers from the point of view of a ‘generalized’ Liouville equation. In particular, given a dynamical system whose phase trajectories are the integral curves of a vector field $X$, the equation (2.2) may be re-written as
$X(\ln|M|)+{\rm div}X=0,$ (5.1)
where we have taken the modulus sign in writing $|M|$ as is necessary for interpreting it as a phase-space density. The above-mentioned equation (5.1) may be straightforwardly interpreted as a generalized Liouville equation (see also, [27]) where $|M|$ can be viewed as the phase-space density. For conservative mechanical systems described by Hamiltonian dynamics on symplectic manifolds (where $X=X_{H}$; as defined in (2.6)), from £$\pounds_{X_{H}}\omega^{n}=0$, one has ${\rm div}X_{H}=0$ which gives $X_{H}(\ln|M|)=0=X_{H}(|M|)$. This is the Liouville equation that one encounters in classical mechanics and statistical mechanics, wherein one identifies the phase-space density $\rho\equiv|M|$. Explicitly, one has
$\frac{d\rho}{dt}=0,$ (5.2)
which is foundational to the formulation of statistical mechanics. In the present study, we have considered systems that cannot be described from the point of view of Hamiltonian dynamics on symplectic manifolds and have presented explicit results for invariant phase-space measures $\{\Omega_{\rm Invariant}=M\Omega\}$ starting with volume-forms $\{\Omega\}$ that are not conserved under the dynamics, i.e., £$\pounds_{X}\Omega\neq 0$. Different cases for $X$ as discussed in this work are conformal ($X^{\gamma}_{H}$), generalized conformal ($X^{K}_{H}$), and contact ($X_{h}$) vector fields. Thus, the present exposition can be understood as a preliminary step towards a formulation of statistical mechanics for dissipative systems (see for example, [11, 13]).